OpenStudy (anonymous):

Heelpp plzz one number is 3 more than another. the sum of their squares is 89 ,what are the numbers?

OpenStudy (anonymous):

let first no is x....then other x+3.. (x)^2 +(x+3)^2 = 89 now solve for x

OpenStudy (mathteacher1729):

One number is three more than another... So we're dealing with two numbers, N and (N+3). The sum of their squares... (N)^2 plus (N+3)^2 ... is 89 so... (N)^2 + (N+3)^2 = 89 What are the numbers? Solve for N. That gives you the first number, then add three to it to find the other. Make sure to check to make sure that the sum of their squares really does add up to 89.

OpenStudy (anonymous):

5 and 8

OpenStudy (mathteacher1729):

So there are two phases to this problem: 1) the setup (obtaining the equation) N^2 + (N+3)^2 = 89 and 2) solving the equation Usually students have the hardest time with step (1).

OpenStudy (earthcitizen):

when x+3=y http://www.wolframalpha.com/input/?i=x%5E2%2By%5E2%3D89

OpenStudy (earthcitizen):

x = ±8, y = ±5

OpenStudy (earthcitizen):

x = ±5, y = ±8

OpenStudy (anonymous):

yup